# Deep Optimal Transport for Domain Adaptation on SPD Manifolds

### v4

@@ III-B Deep Optimal Transport
Overall Objective Function
The overall objective function of DOT is composed of the cross-entropy loss $\mathcal{L}_{CE}$ and the joint distribution adaptation, i.e., $\mathcal{L}_{DOT}:=\alpha_{1}\mathcal{L}_{CE}+\alpha_{2}\mathcal{L}_{MDA}^{2}+\alpha_{3}\mathcal{L}_{CDA}^{2}$, where $\alpha_{1},\alpha_{2}$, and $\alpha_{3}\geq 0$.
Remark.
1). $\mathcal{L}_{MDA}$ and $\mathcal{L}_{CDA}$ should be in the squared form, as required by Brenier’s formulation and the associated theoretical results [32].
[…]
3). In practice, we substitute Equation 1 into $\mathcal{L}_{MDA}$ and $\mathcal{L}_{CDA}$, resulting in the simplified expressions,
$\mathcal{L}_{MDA}$ $=||\sum_{S_{i}\in\mathcal{B}_{S}}\frac{\log{(S_{i})}}{|\mathcal{B}_{S}|}-\sum_{S_{j}\in\mathcal{B}_{T}}\frac{\log{(S_{j})}}{|\mathcal{B}_{T}|}||_{\mathcal{F}};$
$\mathcal{L}_{CDA}$ $=||\sum_{{S^{\ell}_{i}}\in\mathcal{B}_{S^{l}}}\frac{\log{({S^{\ell}_{i}})}}{|\mathcal{B}_{S^{\ell}}|}-\sum_{{S^{\ell}_{j}}\in\mathcal{B}_{T^{\ell}}}\frac{\log{({S^{\ell}_{j}})}}{|\mathcal{B}_{T^{\ell}}|}||_{\mathcal{F}}.$
These simplified expressions have two distinct interpretations in domain adaptation as follows: First, they can be seen as an empirical maximum mean discrepancy approach defined as, $\mathcal{L}_{MMD}=||\sum_{i=1}^{n}\varphi{(S_{i})}/n-\sum_{j=1}^{m}\varphi{(S_{i})}/m||_{\mathcal{H}}$, where $n=|\mathcal{B}_{S}|$, $m=|\mathcal{B}_{T}|$, $\varphi{(x)}:=\log{(x)}$ is the feature map, $\mathcal{H}:=\mathbb{R}^{d}$ is the reproducing kernel Hilbert space, and $d$ is the output dimension; Second, since each matrix $S$ in $\mathcal{B}_{S}$ or $\mathcal{B}_{T}$ is a covariance matrix of multichannel signals, the expressions can also be viewed as a correlation alignment (CORAL) approach [33, 34]. This approach aligns the second-order statistics of the source and target data distributions to minimize the drift between their statistical distributions.
4). For multi-source adaptation, the loss function can be modified by incorporating multi-joint distribution adaptions between each pair of source and target domains on ($\mathcal{S}_{++}$, $g^{LEM}$), as $\mathcal{L}:=\alpha_{1}^{i}\mathcal{L}_{CE}+\sum_{i}\Big{(}\alpha_{2}^{i}\mathcal{L}_{MDA,i}+\alpha_{3}^{i}\mathcal{L}_{CDA,i}\Big{)}$, where the subscript $i$ is used to represent the $i^{th}$ source domain. This modification is based on the theory of multi-marginal optimal transport on Riemannian manifolds [35] and multi-source joint distribution optimal transport [36].
5). DOT does not directly solve an OT problem. Instead, it uses the transportation costs as a guiding principle to direct the learning process of the neural network. Consequently, the trained neural network maps the source and target domains onto a submanifold of the SPD manifold, thereby implementing a subspace method in transfer learning.
6). DOT calculates this cost for the centroid points and addresses the issue of conditional marginal adaptation without necessitating alternating parameter updates. In contrast, DeepJDOT computes the transportation costs for all points and requires alternating parameter updates, as the transport plan is a deterministic transfer.
IV EXPERIMENTS
IV-A Deep Optimal Transport-Based Motor Imagery Classifier
We model the space of EEG spatial covariance matrices in ($\mathcal{S}_{++}$, $g^{LEM}$) and employ DOT to address the issue that EEG statistical characterization is shifted during the feedback phase. Using the event-related desynchronization/synchronization methods to classify motor imagery classes, EEG frequency components in specific frequency bands are found which provided the best discrimination between movement imagination [37]. Hence, it is reasonable to regard features in source and target feature space holds the same, i.e., $X_{S}=X_{T}$, after transforming to a latent space using the Bi-Map layer, and thus model the space of EEG spatial covariance matrices using SPD manifolds with the Log-Euclidean metric. Moreover, the Bi-Map layer provides a transitive group action on SPD manifolds, allowing any transition between two points on manifolds. In Fig. 2, we illustrate how the DOT-based MI-EEG classifier functions in a domain adaptation problem.
Specifically, EEGs are represented by $X\in\mathbb{R}^{n_{C}\times n_{T}}$, where $n_{C}$ and $n_{T}$ denote the number of channels and timestamps, respectively. The corresponding EEG spatial covariance matrix, $S:=S(\Delta f\times\Delta t)\in\mathcal{S}_{++}^{n_{C}}$, represents an EEG segment within the frequency band $\Delta f$ and time interval $\Delta t$. Let $\mathcal{F}$ and $\mathcal{T}$ represent the frequency and time domains, respectively. A given segmentation $Seg(\mathcal{F},\mathcal{T})$ is written as $\{\Delta f\times\Delta t|\Delta f\subset Range(\mathcal{F})\text{, and }\Delta t\subset Range(\mathcal{T})\}$, on the time-frequency domain of EEG signals. To ensure the theoretical convergence of this classifier in a real-world context, the following three assumptions are necessary:
Assumption 1.
There exist linear transformations between the feature space distributions on the source and target domains for each segment of $Seg(\mathcal{F},\mathcal{T})$, which can be estimated using the Kantorovich Problem 2.
Assumption 2.
[…]
Assumption 3.
@@ IV-A Deep Optimal Transport-Based Motor Imagery Classifier
where $\mathcal{B}_{T}^{\prime}$ is a batch of samples on the target domain that is generated by the other time-frequency segments $\Delta f^{\prime}\times\Delta t^{\prime}\in Seg(\mathcal{F},\mathcal{T})$, where $\Delta f^{\prime}\neq\Delta f$ or $\Delta t^{\prime}\neq\Delta t$.
[…]
Assumption 1 is a necessary condition of modeling the OT-DA problem. Assumption 2 is straightforward because when we are uncertain about which EEG frequency band has a higher weight for a specific task, we tend to assume that each frequency band has equal priority for the classification task. Assumption 3 ensures that features extracted from different bands do not overlap.
@@
Suppose EEG signals in the source and target domain have the same segmentation, Seg($\mathcal{F},\mathcal{T}$). Assuming we model EEG spatial covariance matrices in ($\mathcal{S}_{++}^{n_{C}}$, $g^{LEM}$) and that Assumption 1,  2, and 3 exist, then the values of the continuous optimal transport $T(S)$ reside within the set $\big{\{}\log(S_{i})\big{\}}_{i=1}^{N}$.
@@ IV-B Evaluation Dataset
Three commonly used EEG-based motor imagery datasets are used for evaluation, accessed from two publicly available data packages MOABB and BNCI Horizon 2020111  The MOABB package can be accessed at https://github.com/NeuroTechX/moabb, and the BNCI Horizon 2020 project is available at http://bnci-horizon-2020.eu/database/data-sets.. The selection criteria require each subject in the dataset to have two-session recordings collected on different days. All digital signals were filtered using Chebyshev Type II filters at intervals of 4 Hz. The filters were designed to have a maximum loss of 3 dB in the passband and a minimum attenuation of 30 dB in the stopband. The following are the descriptions of these three datasets.
@@
1). Korea University (KU) Dataset: The KU dataset comprises EEG signals obtained from 54 subjects during a binary-class imagery task. The EEG signals were recorded using 62 electrodes at a sampling rate of 1,000 Hz. For evaluation, we selected 20 electrodes located in the motor cortex region. The dataset was divided into two sessions, S1 and S2, each consisting of a training and a testing phase. Each phase included 100 trials, with an equal number of right and left-hand imagery tasks. The EEG signals were epoched from second 1 to second 3.5 relative to the stimulus onset, resulting in a total epoch duration of 2.5 seconds.
@@
2). BNCI2014001 Dataset: The BNCI2014001 dataset (also known as BCIC-IV-2a) involved 9 participants performing a motor imagery task with four classes: left hand, right hand, feet, and tongue. The task was recorded using 22 Ag/AgCl electrodes and three EOG channels at a sampling rate of 250 Hz. The recorded data were filtered using a bandpass filter between 0.5 and 100 Hz, with a 50 Hz notch filter applied. The study was conducted over two sessions on separate days, with each participant completing a total of 288 trials (six runs of 12 cue-based trials for each class). The EEG signals were epoched from the cue onset at second 2 to the end of the motor imagery period at second 6, resulting in a total epoch duration of 4 seconds.
@@
3). BNCI2015001 Dataset: BNCI2015001 involves 12 participants performing a right-hand versus both-feet imagery task. Most of the study consisted of two sessions conducted on consecutive days, with each participant completing 100 trials for each class, resulting in a total of 200 trials per participant. Four subjects participated in a third session, but their data should have been included in the evaluation. The EEG data were acquired at a sampling rate of 512 Hz and filtered using a bandpass filter ranging from 0.5 to 100 Hz, with a notch filter at 50 Hz. The recordings began at second 3 with the prompt and continued until the end of the cross period at second 8, resulting in a total duration of 5 seconds.
IV-C Experimental Settings
@@
The experimental settings for domain adaptation on three evaluated datasets are described as follows, illustrated in Fig. 3: 1). Semi-supervised Domain Adaptation: For the KU dataset, S1 was used for training, and the first half of S2 was used for validation, while the second half was reserved for testing. Early stopping was adopted during the validation phase. 2). Unsupervised Domain Adaptation: For the BNCI2014001 and BNCI2015001 datasets, S1 was used for training, and S2 was used for testing. No validation set was used, and a maximum epoch value was set for stopping in practice.
We specified that the transfer procedure must always occur from S1 (source domain) to S2 (target domain), i.e., $S1\rightarrow S2$ for each dataset. It is reasonable because there is no meaning if negative transfer is used in the learning process. The former and latter sessions could be interpreted as the calibration and feedback phases.
@@
IV-D Evaluated Baselines
The geometric methods in this study can be broadly divided into four categories. These include categories of parallel transport, deep parallel transport, optimal transport, and deep optimal transport. Each category includes various transfer methods, performing on three types of base machine learning classifiers: the support vector machine, minimum distance to Riemannian mean, and SPDNet. We begin by introducing these base classifiers and discussing transfer methods in each category.
@@ IV-D Evaluated Baselines
    Minimum Distance to Riemannian Mean (MDM) [21]: This employs the geodesic distance on SPD manifolds for classification. More specifically, a centroid is estimated for each predefined class, and subsequently, the classification of a new data point is determined based on its proximity to the nearest centroid.
@@
    SPDNet [29]: This has been introduced in the preliminary part. In the experiments, we employ a one-depth Bi-Map layer with both the input and output dimensions equal to the number of electrodes on the primary motor cortex: 20 (KU), 22 (BNCI2014001), and 13 (BNCI2015001).
As computations on the SPD manifold are related to Riemannian metrics, we will emphasize the specific Riemannian metrics associated with each method. The different transfer methods across categories according to various Riemannian metrics can be seen in the first three columns of Table II.
IV-D1 Category of Parallel Transport
[…]
IV-D2 Category of Deep Parallel Transport
This category refers to the domain adaptation approaches using deep learning and parallel transport on SPD manifolds. One representative work is known as Riemannian Batch Normalization (RieBN) [41]. RieBN is an SPD matrix-valued network architecture that employs parallel transport for batch centering and biasing on SPD neural networks.
IV-D3 Category of Optimal Transport
This category refers to treating the distribution of the covariance matrix obtained from cross-session as an OT problem and solving the transport plan using different variants of Wasserstein distances [42]. The classifier is trained within a session and tested in a shifted session. This is an unsupervised method. Three different types of Wasserstein distances are provided, including the earth mover’s distance (EMD), Sliced-Wasserstein distance (SPDSW), and Log Sliced-Wasserstein distance (logSW).
IV-D4 Category of Deep Optimal Transport
This category refers to the approach proposed in this study, which uses SPD neural networks on ($\mathcal{S}_{++}$, $g^{LEM}$) to transfer both source domain and target domain to a common space induced by optimal transport.
V Main Results
Figure 4: Synthetic Data of 2-dimensional SPD Cone: The subfigures from left to right and top to bottom are labeled as (a) to (d). In subfigures (a) to (d), the red points represent the source domain, while the blue points indicate their locations after the transformation.
TABLE II: Experimental Results: All results in the table are in percentages (%), and the comparison methods include four categories, nine methods, and the corresponding three base classifiers. The original results of the base classifiers without considering transfer methods are at the bottom of the table. In the table, we use the symbol "/" to indicate that this method does not exist in the scenario constructed by this dataset.
Metric Category Method Classifier KU (54 Subjects) BNCI2014001 (9 Subjects) BNCI2015001 (12 Subjects)
@@ V Main Results
Semi-supervised Unsupervised Unsupervised
Accuracy Increment Accuracy Increment Accuracy Increment
@@
V-A Synthetic Dataset
We synthesize a source domain dataset consisting of 50 covariance matrices of size $2\times 2$, drawing a Gaussian distribution on the SPD manifold proposed in a generative approach [43] with a median of 1 and a standard deviation of 0.4. We utilize the following transformation matrix in the generation,
$W:=\begin{pmatrix}1&0.5\\ 0.5&1\\ \end{pmatrix},$
and data in the target domain can be expressed as $WSW^{\top}$.
Subfigure (a) exhibits red points sourced from the source domain, with the blue points originating from the target domain. In Subfigure (b), the blue points represent the relocation of the red points via the EMD method from OT. The associations made by this transformation are represented by thin blue lines. Remarkably, the position of these transformed blue points is extremely proximate to the points of the target domain in Subfigure (a). This can be attributed to the fact that the OT-DA method essentially computes a weighted average of the source and target domain data points, based on weights learned to maximize the transportation costs.
The transformed blue points in Subfigures (c) and (d) were derived from mapping the red points via neural network weights learned through the DOT and DeepJDOT methods, respectively. Corresponding point pairs are linked with thin blue lines. These figures reveal the shared fundamental characteristic of the DOT and DeepJDOT methods that both project the source and target domains onto a subspace learned via a network, which, in these two figures, is represented by a straight line. In differential geometry, this is a submanifold, which refers to a lower-dimensional manifold embedded within a higher-dimensional one. In this case, the two-dimensional matrices are represented by a line, essentially projecting the higher dimensional data onto a lower dimensional space. It reduces the complexity of the data while preserving information.
V-B Comparative Experiments on Real-World EEG Datasets
As shown in Table II, we classify all approaches into three categories according to Riemannian metrics. The first category does not require a Riemannian metric for computation, the second employs $g^{AIRM}$, while the third uses $g^{LEM}$.
The experimental results indicate that metrics strongly influence the results. Specifically, in the table, EMD+SVM achieved 50.89 ($g^{AIRM}$) and 68.33 ($g^{LEM}$) on the BNCI2014001 dataset. Moreover, several observations are introduced as follows: 1). Participant Number: The improvement is related to the number of participants. On datasets with fewer participants, the enhancement brought by transfer methods is more significant, with several scenarios showing changes of more than 5%. 2). Base Classifier Results: The extent of improvement is linked to the base classifier’s performance. Lower initial results from the base classifier provide more opportunities for transfer methods to achieve significant gains. 3). Methodological Factors: Improvements are associated with the classifier neural network architecture, parameter selection, and data preprocessing. Therefore, the conclusions of this study may not be universally applicable.
Furthermore, we summarize the observations of all methods affected by various Riemannian metrics as follows:
  • ($\mathcal{S}_{++}$, $g^{AIRM}$) includes categories of OT, PT, and DPT. Combining the results from all three datasets, the RCT method performs well, especially when paired with the MDM or SPDNet classifier. The EMD method shows the largest variation across datasets. The second step (ROT) in the RPA method, which is limited to semi-supervised scenarios, performs worse than the first step (RCT) of RPA. The RieBN method provides slight improvements but does not outperform RCT across the board.
  • ($\mathcal{S}_{++}$, $g^{LEM}$) includes two categories of OT and DOT. The EMD method, when paired with the SVM classifier, performs significantly better compared to its pairing with the SPDNet classifier. SPDSW and logSW methods have obvious improvements compared to their base classifier (SVM), among which SPDSW performs better than logSW by 1% to 2% on all three datasets. The MDA and CDA methods perform similarly, except for the BNCI2014001 dataset, the improvements on the other two datasets are between 1% to 2%, while the MDA method is slightly better than the CDA method or a mixture of the MDA and CDA methods. Notably, the use of pseudo-labels in the CDA method affects its results. The DeepJDOT method’s performance across three datasets falls short of our proposed approach. The primary reason for this could be attributed to our decision to replace the task of computing transportation costs for each data point with merely computing transportation costs for the centers of each frequency band’s dataset. Minimizing the transportation costs does not necessarily imply improved classification accuracy, thus the proposed approach reduces computational cost while enhancing the model’s potential.
Besides, neural network-based approaches generally achieve better classification outcomes compared to other categories of approaches, such as EMD, SPDSW, and logSW. To compute the CDA loss in DOT, we used the pseudo-label method, with pseudo labels derived from the Graph-CSPNet [26], which currently performs the best for subject-specific tasks. Specifically, the results of SPDSW and logSW on the BNCI2014001 dataset were obtained using the data processing procedure and source codes provided by their GitHub repository, outperforming ours. This is because the performance of the SVM methods heavily depends on a selection of handcrafted features.
VI DISCUSSIONS
The experimental results indicate that all transfer learning methods exhibit effectiveness. However, it’s important to note that while domain shift is a known factor in these tasks, the classification performance is not solely influenced by it. Consequently, the improvements from transfer learning methods are generally limited. Currently, we lack an effective method to quantify the intensity of domain shift and a system to correlate this intensity with classifier performance, making it difficult to fully assess these methods’ effectiveness. In the rest of this section, we will discuss two issues of the proposed approach.
VI-A Theorem 3 with cost function $c(x,\bar{x})=||x-\bar{x}||_{\ell^{2}}^{2}$

### v5

@@ 4.1 Deep Optimal Transport-Based Motor Imagery Classifier
Assumption 3 (Feature Disjointness).
@@
where $\mathcal{B}_{T}^{\prime}$ is a batch of target samples obtained from a different segment $\Delta f^{\prime}\times\Delta t^{\prime}\in\text{segment}(\mathcal{F},\mathcal{T})$, such that $\Delta f^{\prime}\neq\Delta f$ or $\Delta t^{\prime}\neq\Delta t$.
Figure 2: Illustration of DOT-Based Motor-Imagery Classifier: The proposed neural networks $\varphi$ simultaneously transports the Fréchet means $\overline{w}(\mathcal{B}_{S})$ and $\overline{w}(\mathcal{B}_{T}$ of each batch of source and target samples within each frequency band to a common subspace $Z$ associated with that frequency band. Assumption 2 guarantees that all frequency components are equally weighted in the classification process. Assumption 3 ensures that the transformations applied to the source and target domains are restricted to occur within corresponding EEG frequency bands.
Remark.
1). Assumption 1 establishes the foundational requirements for formulating the OT-DA problem on Riemannian manifolds.
2). To address uncertainty in EEG frequency band weighting for motor imagery tasks, Assumption 2 imposes a uniform prior across all spectral components. This implies equal contribution weights for all frequency bands in the classification objective, unless task-specific evidence dictates otherwise.
3). Assumption 3 enforces non-overlapping feature distributions across distinct frequency bands, ensuring orthogonality in the latent representations learned from different spectral regions.
@@
Suppose EEG signals in the source and target domain have the same segmentation, segment($\mathcal{F},\mathcal{T}$). Assuming we model EEG spatial covariance matrices in ($\mathcal{S}_{++}^{n_{C}}$, $g^{LEM}$) and that Assumption 1,  2, and 3 exist, then the values of the continuous optimal transport $T(S)$ reside in the set $\big{\{}\log(S_{i})\big{\}}_{i=1}^{N}$.
@@ 4.2 Evaluation Dataset
We evaluate our method on three publicly available EEG motor imagery datasets—Korea University (KU), BNCI2014001 (BCIC-IV-2a), and BNCI2015001—accessed via the MOABB package (https://github.com/NeuroTechX/moabb) and BNCI Horizon 2020 project (http://bnci-horizon-2020.eu/database/data-sets). All datasets include two-session recordings per subject (collected on different days) and were preprocessed using Chebyshev Type II filters with a 4 Hz bandwidth, 3 dB passband ripple, and 30 dB stopband attenuation. The descriptions of the three datasets are as follows:
@@
1). Korea University (KU) Dataset: Also referred to as Lee2019_MI, this dataset contains EEG recordings from 54 subjects performing a binary-class motor imagery task. EEG signals were acquired using 62 electrodes at a sampling rate of 1,000 Hz. For analysis, we focused on 20 electrodes located over the motor cortex. Each subject participated in two sessions (S1 and S2), each consisting of training and testing phases with 100 trials per phase, equally split between left- and right-hand imagery tasks. EEG epochs were extracted from 1 to 3.5 seconds post-stimulus, yielding a segment duration of 2.5 seconds.
@@
2). BNCI2014001 Dataset: Also known as BCIC-IV-2a, this dataset comprises EEG signals from 9 subjects performing a four-class MI task (left hand, right hand, feet, tongue). EEG was recorded from 22 Ag/AgCl electrodes and three EOG channels at a sampling rate of 250 Hz. Signals were bandpass filtered between 0.5$\sim$100 Hz and notch filtered at 50 Hz. Each subject completed two sessions on different days, with 288 trials per subject, divided into six runs of 12 cue-based trials per class. Epochs were segmented from 2 to 6 seconds post-cue, resulting in a 4-second window.
@@
3). BNCI2015001 Dataset: This dataset features 12 participants engaged in a binary MI task (right hand vs. both feet). Most participants completed two sessions on consecutive days, each comprising 100 trials per class (200 trials in total). Four participants also completed a third session, which was included in our evaluation. EEG was recorded at a sampling rate of 512 Hz and bandpass filtered between 0.5$\sim$100 Hz with a 50 Hz notch filter. Epochs were extracted from 3 to 8 seconds post-prompt, yielding a 5-second duration.
4.3 Experimental Settings
@@
For domain adaptation, experiments followed two protocols, illustrated in Fig. 3: 1). Semi-supervised Domain Adaptation: For the KU dataset, S1 was used for training, and the first half of S2 was used for validation, while the second half was reserved for testing. Early stopping was adopted during the validation phase. 2). Unsupervised Domain Adaptation: For the BNCI2014001 and BNCI2015001 datasets, S1 was used for training, and S2 was used for testing. No validation set was used, and a maximum epoch value was set for stopping in practice.
We specified that the transfer is strictly unidirectional, i.e., from the source domain to the target domain $S1\rightarrow S2$, reflecting real-world calibration-to-feedback workflows and avoiding counterproductive negative transfer.
@@
4.4 Evaluated Baselines
The geometric methods employed in this study can be broadly categorized into four groups: parallel transport, deep parallel transport, optimal transport, and deep optimal transport. Each category encompasses distinct transport methodologies that are systematically applied to three fundamental machine learning classifiers: Support Vector Machines (SVM), Minimum Distance to Riemannian Mean (MDRM), and SPDNet. We first present an overview of these base classifiers before detailing the transport methods within each methodological category.
@@ 4.4 Evaluated Baselines
    Minimum Distance to Riemannian Mean (MDRM) [10]: This classifier employs the geodesic distance on SPD manifolds for classification. More specifically, a centroid is estimated for each predefined class, and subsequently, the classification of a new data point is determined based on its proximity to the nearest centroid.
@@
    SPDNet [27]: This classifier was described in the previous section. In the experiments, we utilize a single-layer BiMap architecture where the input and output dimensions align with the number of electrodes in the primary motor cortex: 20 (KU dataset), 22 (BNCI2014001 dataset), and 13 (BNCI2015001 dataset).
Since all computations on the SPD manifold inherently involve Riemannian metrics, we explicitly identify the Riemannian metric associated with each method. The variations in transport methods across methodological categories, distinguished by their underlying Riemannian metrics, are systematically organized in the first three columns of Table 2.
  1. i.
    Category of Parallel Transport (PT): This category encompasses methods that apply parallel transport to align training and test data distributions across domains. A prominent technique in this framework is Riemannian Procrustes Analysis (RPA) [40], which operates on the SPD manifold equipped with the Affine-Invariant Riemannian Metric. RPA employs three sequential geometric transformations:
    1. (1).
      Recentering (RCT): Aligns dataset centroids to the identity matrix via parallel transport from the Fréchet mean (of either source or target domain) to $I$.
    2. (2).
      Stretching: Rescales data distributions to equalize dispersion around the mean.
    3. (3).
      Rotation (ROT): Adjusts the orientation of the target domain’s point cloud to match the source domain.
    Notably, ROT functions as an unsupervised domain adaptation step, while RCT is restricted to semi-supervised settings due to its dependency on labeled source data. The RCT method, originally proposed as an independent solution for cross-session motor imagery classification [41], is integrated here as the foundational alignment step in RPA. Post-transformation, classification is performed using the MDRM algorithm.
  2. ii.
    Category of Deep Parallel Transport (DPT): This framework integrates deep learning with parallel transport operations on the SPD manifold. A key implementation is Riemannian Batch Normalization (RieBN) [42], an SPD-valued neural network architecture that leverages parallel transport for batch normalization. RieBN performs intrinsic centering and bias correction directly on the manifold, ensuring geometrically consistent feature distributions across layers.
  3. iii.
    Category of Optimal Transport (OT): These methods formulate cross-session covariance matrix distribution alignment as an OT problem, solved using Wasserstein distance variants. The classifier is trained on session-specific data and evaluated under session-shifted conditions, operating in a fully unsupervised regime. Three Wasserstein metrics [43] are applied:
    1. (1).
      Earth Mover’s Distance (EMD);
    2. (2).
      Sliced Wasserstein Distance on SPD matrices (SPDSW);
    3. (3).
      Log-Euclidean Sliced Wasserstein Distance (logSW).
  4. iv.
    Category of Deep Optimal Transport (DOT): The proposed method unifies SPDNet and OT on the SPD manifold equipped with Log-Euclidean Metric. By jointly optimizing a transport plan and a deep feature extractor, it projects source and target domains into an OT-aligned latent space, enabling invariant classification under domain shifts.
5 Main Results
Figure 4: Comparison of OT and DOT on 2 by 2 SPD matrices: Each colored ball represents a 2 by 2 SPD matrix. The left diagram illustrates the OT method, where the source distribution (red) is mapped to the target distribution (green), with blue lines indicating the transport paths between each pair. The right diagram depicts the proposed DOT approach but only considers the marginal distribution adaptation, where both the source and target distributions are mapped onto the subspaces (brown) respectively.
5.1 Synthetic Dataset
We synthesize a source domain dataset consisting of SPD matrices of size $2\times 2$, drawing a Gaussian distribution on the SPD manifold proposed in a generative approach [44] with a median of 1 and a standard deviation of 0.4. We adopt the following transformation matrix in the generation:
$W:=\begin{pmatrix}1&0.5\\ 0.5&1\\ \end{pmatrix},$
and data in the target distribution can be expressed as $WSW^{\top}$.
This experiment does not generate label information, so the DOT method we use only considers marginal distribution adaptation. From Fig. 4, we observe that the DOT method effectively learns a mapping $\varphi$ that aligns the source and target distributions $X_{S}$ and $X_{T}$ onto the subspaces $\varphi(X_{S})=\varphi(X_{T})$ by minimizing the difference between their means after mapping.
5.2 Comparative Experiments on Real-World EEG Datasets
As summarized in Table 2, we categorize all evaluated methods into three groups based on their reliance on Riemannian metrics: (1) non-metric approaches, (2) methods using the Affine-Invariant Riemannian Metric, and (3) those employing the Log-Euclidean Metric. Results demonstrate significant metric-dependent performance variations. For instance, on the BNCI2014001 dataset, EMD+SVM achieves 50.89% accuracy under $g^{AIRM}$ versus 68.33% under $g^{LEM}$.
Key empirical insights emerge: (i) Participant scale: Transfer methods yield more pronounced improvements (often >5%) in smaller cohorts due to heightened inter-subject variability; (ii) Baseline dependency: Methods achieve larger gains when base classifiers (e.g., SVM, MDRM, SPDNet) underperform initially; (iii) Methodological sensitivity: Performance correlates with architectural choices (e.g., neural network design), hyperparameter tuning, and preprocessing pipelines, limiting universal generalizability.
Table 2: Experimental Results: All results in the table are expressed in percentages (%) and include comparisons across four categories, nine methods, and their corresponding three base classifiers. The baseline results of these classifiers (base classifier), prior to the application of any transfer methods, are presented at the bottom of the table. The abbreviation "Acc." stands for average accuracy, and the symbol $\uparrow$ highlights the performance improvements achieved by incorporating transfer methods compared to the baseline results without them.
Metric Category Method Base KU BNCI2014001 BNCI2015001
(54 Subjects) (9 Subjects) (12 Subjects)
@@ 5.2 Comparative Experiments on Real-World EEG Datasets
Classifier Semi-supervised Unsupervised Unsupervised
Acc. $\uparrow$ Acc. $\uparrow$ Acc. $\uparrow$
@@
For methods based on the affine-invariant Riemannian metric, the recentering technique (RCT) consistently achieves strong performance across datasets, especially when combined with MDRM or SPDNet classifiers. In contrast, EMD exhibits high variability across datasets, indicating limited robustness. Within the RPA framework, the rotation step (ROT) underperforms its recentering counterpart and is further constrained by its semi-supervised nature. RieBN yields only marginal improvements and fails to consistently outperform RCT.
Under the Log-Euclidean metric, the performance of OT-based methods diverges notably. EMD combined with SVM surpasses its combination with SPDNet, while SPDSW and logSW outperform the base SVM by 1$\sim$2%, with SPDSW consistently achieving the best results. MDA shows slight advantages over CDA on two datasets (1$\sim$2improvement), possibly due to CDA’s reliance on unstable pseudo-labels generated by Graph-CSPNet [24]. In particular, DeepJDOT performs poorly on our proposed method. Although our approach simplifies alignment by minimizing transport costs between frequency-band centroids, trading off direct accuracy for computational efficiency, it significantly improves scalability without compromising robustness.
Overall, neural network-based approaches consistently outperform traditional methods such as EMD and SPDSW. Although feature engineering remains critical for SVM-based techniques, deep learning models benefit from the representations learned. Although public implementations of SPDSW and logSW yield better results than ours on BNCI2014001 - likely due to task-specific preprocessing - their dependence on handcrafted features highlights a key limitation compared to the flexibility and generalizability of deep models.
6 DISCUSSIONS
In the following discussion, we address several critical limitations inherent to our proposed methodology:
6.1 Theorem 3 with cost function $c(x,\bar{x})=||x-\bar{x}||_{\ell^{2}}^{2}$